a garden store sells bermuda grass seed for $3.65 per pound and argentine bahia grass seed for $6.30 per pound. how many pounds of bermuda grass seed should be mixed with 20 pounds of bahia grass seed to obtain a low maintenance and drought resistance blend that will sell for $4.90?
Answer by Alan3354(69443)

Same problem, different numbers
=======================================================
how many liters of a 30% salt solution should be mixed with a 80% salt solution to obtain 120 liters of a 60% salt solution?
-----------
t = amount of 30%
e = amount of 80% --- good luck finding 80% salt
----
t + e = 120 (total solution)
30t + 80e = 120*60 ( total salt)
etc